# Sparse Actuation for LPV Systems with Full-State Feedback in $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ Framework

### v1

@@ III New Contributions
IV Problem Formulation
IV-A Brief Overview of LPV Systems
@@
IV-B Sparse Actuation for Full State-Feedback LPV Systems
@@
where $\text{row}_{i}(K)\in\mathcal{R}^{1\times N_{x}}$ is the $i^{th}$ row of the controller $K$. A smaller value of $\|\mathcal{G}_{u_{i}}(\rho,s)\|_{\mathcal{H}_{2}}$ indicates that the contribution of the $i^{th}$ actuator to the peak magitude of control is small. If this norm value falls below a certain threshold (e.g., 1E-3), it implies that the $i^{th}$ actuator may be redundant and could be removed from the control architecture. In essence, a vector with sparse $\mathcal{H}_{2}$ norm values from $w$ to $u_{i}$ corresponds to an actuation structure with fewer active actuators. This is the basis for the sparse controller design approach discussed below.
@@
IV-C Iterative Reweighted $l_{1}$ Minimization
@@ IV-D System Model
The resulting LPV system has $N_{x}$ parameters and is affine in nature. The physical parameters of the system are given in Table I.
@@ V Results
V-A Sparse Actuation with $\mathcal{H}_{\infty}$ Full-State Feedback
@@
Figure 2: Angular deflection trajectories from nonlinear simulations for a given $\mathcal{H}_{\infty}$ performance.
Figure 3: Angular deflection rate trajectories from nonlinear simulations for a given $\mathcal{H}_{\infty}$ performance.
@@
V-B Sparse Actuation with $\mathcal{H}_{2}$ Full-State Feedback
@@
Figure 6: Angular deflection trajectories from nonlinear simulations for a given $\mathcal{H}_{2}$ performance.
Figure 7: Angular deflection rate trajectories from nonlinear simulations for a given $\mathcal{H}_{2}$ performance.

### v2

@@ II New Contributions
III Problem Formulation
III-A Brief Overview of LPV Systems
@@
III-B Sparse Actuation for Full State-Feedback LPV Systems
@@
where $\text{row}_{i}(K)\in\mathcal{R}^{1\times N_{x}}$ is the $i^{th}$ row of the controller $K$. A smaller value of $\|\mathcal{G}_{u_{i}}(\rho,s)\|_{\mathcal{H}_{2}}$ indicates that the contribution of the $i^{th}$ actuator to the peak magitude of control is small. If this norm value falls below a certain threshold (e.g., $10^{-3}$), it implies that the $i^{th}$ actuator may be redundant and could be removed from the control architecture. In essence, a vector with sparse $\mathcal{H}_{2}$ norm values from $w$ to $u_{i}$ corresponds to an actuation structure with fewer active actuators. This is the basis for the sparse controller design approach discussed below.
@@
III-C Iterative Reweighted $l_{1}$ Minimization
@@ III-D System Model
The EOM for this problem is nonlinear as the states appear in quadratic form, which allows for quasi-LPV representation. The resulting LPV system has $N_{x}$ parameters and is affine in nature. The physical parameters of the system are given in Table I.
@@ IV Results
IV-A Sparse Actuation with $\mathcal{H}_{\infty}$ Full-State Feedback
@@
Figure 2: Angular deflection trajectories from nonlinear simulations for the given $\mathcal{H}_{\infty}$ performance $\gamma_{0}=0.15$.
Figure 3: Angular deflection rate trajectories from nonlinear simulations for the given $\mathcal{H}_{\infty}$ performance $\gamma_{0}=0.15$.
@@
IV-B Sparse Actuation with $\mathcal{H}_{2}$ Full-State Feedback
@@
Figure 6: Angular deflection trajectories from nonlinear simulations for the given $\mathcal{H}_{2}$ performance $\gamma_{0}=0.15$.
Figure 7: Angular deflection rate trajectories from nonlinear simulations for the given $\mathcal{H}_{2}$ performance $\gamma_{0}=0.15$.